OpenStudy (anonymous):

6x-12+2x=3+8x=15 Is this always sometime or never true ?

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

Let's solve the first equation first: 6x-12+2x = 8x-12 =15 => 8x = 27

OpenStudy (anonymous):

Now let's solve the second equation 8x+3=15 =>8x = 12

OpenStudy (anonymous):

So there are no values for x which will satisfy both equations.

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

You're most welcome!!! :) :) :)